The wave function for a harmonic wave on a string is y(x, t) = (0.0060 m) sin((53.8 m-1)x + (304 s-1)t). What is its speed? m/s (b) Find the wavelength of this wave. m Find its frequency. Hz Find its period. s (c) What is the maximum speed of any string segment? m/s
Answer #1


Soth: Art) = (.006 m) Sin ( 53.8 x + 3out) (a) v=dy = I = 0.006x 304 cos (53.8x +3 out) 0.0064 dt V=108240609 (53.8 x + 304t)